OpenStudy (anonymous):

Can you show me step by step how to do 1/2+2/3

OpenStudy (anonymous):

first tell me what the greatest common denominator is

OpenStudy (anonymous):

hint multiply the two denominators

OpenStudy (anonymous):

6 , ok will it be a mxed number right ?

OpenStudy (anonymous):

sry lowest common denominator

OpenStudy (anonymous):

7/6

OpenStudy (anonymous):

1/2 X 3 = 3/6 now 2/3 X 2 = 4/6 add them together.... 7/6 and simplify it 1 1/6

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

ok , i gpt 1 1/6

OpenStudy (anonymous):

*got

OpenStudy (anonymous):

yep